# Video-Language Understanding: A Survey from Model Architecture, Model Training, and Data Perspectives

### v2

@@ 4.3 LLM-Augmented Architecture
Large language models (LLMs) have achieved impressive results in simultaneously tackling multiple NLP tasks. Recent efforts have sought to apply LLMs for video-language understanding to extend its cross-domain adaptation ability to video-language settings (Chen et al., 2023; Li et al., 2023a). These efforts can be categorized into two approaches. The first approach employs LLM as a controller and video-language understanding models as helping tools. The controller will call the specific tool according to the language input instruction. The second approach utilizes LLM as the output generator and seeks to align video pre-trained models to the LLM. For video-language understanding, since the second approach dominates the first one with a long list of recent works (Chen et al., 2023; Li et al., 2023a; Chen et al., 2023; Li et al., 2023d; Zhang et al., 2023b; Maaz et al., 2023), we review them as follows:
@@ 5.1 Pre-training for Video-Language Understanding
Contrastive learning has revealed promising results (Lin et al., 2022; Gao et al., 2021; Xue et al., 2022b; Nguyen et al., 2022; Nguyen and Luu, 2021; Nguyen et al., 2024c, a, 2023a; Wu et al., 2023a, 2024, 2022). MLM has contributed to enhancing VideoQA since the task resembles MLM in predicting the language word given a video-language pair (the question is the language input in videoQA). Compared to these pre-training strategies, MVM does provide performance gain for video-language understanding but its gain is less significant. For more details about pre-training, please refer to (Cheng et al., 2023).
@@ 5.2 Fine-tuning for Video-Language Understanding
Fine-tuning strategies. Normally, all of the model parameters are updated during fine-tuning (Gao et al., 2017; Xu et al., 2019; Anne Hendricks et al., 2017; Nguyen et al., 2023b; Wu et al., 2023b). However, in cases computational resources or training data are limited, only adaptation layers such as low-rank adapters (Pan et al., 2022; Yang et al., 2022a; Nguyen et al., 2024b) or learnable prompt vectors (Ju et al., 2022) are fine-tuned to reduce training cost or prevent overfitting. Such risks also apply for LLM-augmented architectures discussed in Section 4.3, since LLMs exhibit a billion scale of parameters, thus incurring excessively huge cost if full fine-tuning is conducted. For such models, Zhang et al. (2023b) and Li et al. (2023d) design a two-stage instruction tuning strategy which only fine-tunes the semantic translator. The first stage trains the model to generate the textual description based on the combined video and the language instruction, in order to align visual representations extracted by the visual encoder with the language space of LLM. The second stage is often performed on small-scale video-text pairs manually collected by the authors to further tailor the output features of the translator towards the target domains.
@@ 6.2 Label annotation
Automatic generation. Directly taking language transcripts of YouTube videos as textual labels could reduce annotation cost (Miech et al., 2019; Xue et al., 2022a; Zellers et al., 2021). However, these labels have been shown to be grammatically incorrect and temporally misalign with the video content (Tang et al., 2021). Motivated by the success of LLMs, Zhao et al. (2023) train a system consisting of a TimeSformer-L visual encoder and a GPT-2XL decoder to write dense captions for videos. Moreover, Li et al. (2023a) use GPT-4 to generate summaries for movie synopses.
@@ 7 Future Directions
Long-form video-language understanding. Current understanding systems have demonstrated remarkable performance on short video clips lasting several seconds. However, they tend to struggle when switching to long-form videos which last several minutes or hours. To enhance the applicability of these systems, it is essential to enhance their capability of understanding long-form videos. Current approaches mainly feature reducing computational cost through architectures more efficient than Transformer-based ones such as state space models (Yang et al., 2024; Li et al., 2024), which can be considered as linear RNN with specifically designed fixed weights, or compensating sparsely extracted video frames with additional information (Lin et al., 2022). In general, how to effectively model long-form videos and adapt them to the joint context with language deserves more attention.

### v3

@@ 4.3 LLM-Augmented Architecture
Large language models (LLMs) have achieved impressive results in simultaneously tackling multiple NLP tasks. Recent efforts have sought to apply LLMs for video-language understanding to extend its cross-domain adaptation ability to video-language settings (Chen et al., 2023; Li et al., 2023a; Nguyen et al., 2025a). These efforts can be categorized into two approaches. The first approach employs LLM as a controller and video-language understanding models as helping tools. The controller will call the specific tool according to the language input instruction. The second approach utilizes LLM as the output generator and seeks to align video pre-trained models to the LLM. For video-language understanding, since the second approach dominates the first one with a long list of recent works (Chen et al., 2023; Li et al., 2023a; Chen et al., 2023; Li et al., 2023d; Zhang et al., 2023b; Maaz et al., 2023), we review them as follows:
@@ 5.1 Pre-training for Video-Language Understanding
Contrastive learning has revealed promising results (Lin et al., 2022; Gao et al., 2021; Xue et al., 2022b; Nguyen et al., 2022; Nguyen and Luu, 2021; Nguyen et al., 2024d, a, 2023a; Wu et al., 2023a, 2024; Nguyen et al., 2025b, c). MLM has contributed to enhancing VideoQA since the task resembles MLM in predicting the language word given a video-language pair (the question is the language input in videoQA). Compared to these pre-training strategies, MVM does provide performance gain for video-language understanding but its gain is less significant. For more details about pre-training, please refer to (Cheng et al., 2023).
@@ 5.2 Fine-tuning for Video-Language Understanding
Fine-tuning strategies. Normally, all of the model parameters are updated during fine-tuning (Gao et al., 2017; Xu et al., 2019; Anne Hendricks et al., 2017; Nguyen et al., 2023b; Wu et al., 2023b). However, in cases computational resources or training data are limited, only adaptation layers such as low-rank adapters (Pan et al., 2022; Yang et al., 2022a; Nguyen et al., 2024c) or learnable prompt vectors (Ju et al., 2022) are fine-tuned to reduce training cost or prevent overfitting. Such risks also apply for LLM-augmented architectures discussed in Section 4.3, since LLMs exhibit a billion scale of parameters, thus incurring excessively huge cost if full fine-tuning is conducted. For such models, Zhang et al. (2023b) and Li et al. (2023d) design a two-stage instruction tuning strategy which only fine-tunes the semantic translator. The first stage trains the model to generate the textual description based on the combined video and the language instruction, in order to align visual representations extracted by the visual encoder with the language space of LLM. The second stage is often performed on small-scale video-text pairs manually collected by the authors to further tailor the output features of the translator towards the target domains.
@@ 6.2 Label annotation
Automatic generation. Directly taking language transcripts of YouTube videos as textual labels could reduce annotation cost (Miech et al., 2019; Xue et al., 2022a; Zellers et al., 2021). However, these labels have been shown to be grammatically incorrect and temporally misalign with the video content (Tang et al., 2021; Nguyen et al., 2024b). Motivated by the success of LLMs, Zhao et al. (2023) train a system consisting of a TimeSformer-L visual encoder and a GPT-2XL decoder to write dense captions for videos. Moreover, Li et al. (2023a) use GPT-4 to generate summaries for movie synopses.
@@ 7 Future Directions
Long-form video-language understanding. Current understanding systems have demonstrated remarkable performance on short video clips lasting several seconds. However, they tend to struggle when switching to long-form videos which last several minutes or hours. To enhance the applicability of these systems, it is essential to enhance their capability of understanding long-form videos. Current approaches mainly feature reducing computational cost through architectures more efficient than Transformer-based ones such as state space models (Yang et al., 2024; Li et al., 2024; Nguyen et al., 2024e), which can be considered as linear RNN with specifically designed fixed weights, or compensating sparsely extracted video frames with additional information (Lin et al., 2022). In general, how to effectively model long-form videos and adapt them to the joint context with language deserves more attention.
@@ Appendix B Analysis of Video-Language Understanding datasets
Ego-QA (Nguyen et al., 2024e) Egocentric videos Manual VideoQA 18.5K
MAD-QA (Nguyen et al., 2024e) Movie videos Manual VideoQA 19K